thank you so much !

Draw the molecule for the following: 1) CH3−CH−CH=CH2+Br2→|CH3 2) CH2=CH−CH3+Cl2→ 3) CH3|CH3-CH3−C−CH=CH2+HCl→|

Spell out full name of the following compounds please:

aH-C-CH2-CH3bcCH H3C-C-CH3
Answer #1

thank you so much! Draw the molecule for the following: 1) CH3-CH-CH=CH2+Br27|CH3 2) CH2=CH-CH3+C12+ 3) CH3CH3-CH3-C-CH=CH2+H